### v1

# Two-Level Distributed Interference Management for Large-Scale HAPS-Empowered vHetNets

Abstract
Next-generation wireless networks (xG) must provide ubiquitous connectivity while enhancing user experience in both densely populated urban areas and rural regions. To achieve this, a disruptive network architecture is essential, and high altitude platform stations (HAPS) offer a promising solution. By integrating HAPS with terrestrial networks, we can create HAPS-empowered vertical heterogeneous networks (vHetNets), which significantly improve coverage and capacity, as well as support emerging use cases. In HAPS-empowered vHetNets, different tiers can share the same spectrum, forming harmonized spectrum vHetNets that enhance spectral efficiency (SE). However, we face two major challenges: i) co-channel interference in harmonized spectrum vHetNets, and ii) the large-scale nature of the network. To address the first challenge, we adopt a cell-free approach as the underlying network architecture for the HAPS-empowered vHetNet. In this approach, base stations use beamforming to direct high-gain, narrow beams toward users, which helps mitigate interference. However, this creates a nonconvex and high-dimensional optimization problem, which highlights the second challenge of dealing with a large-scale network. Consequently, centralized solutions become impractical due to the computational and communication overhead involved. The standard two-block alternating direction method of multipliers (ADMM) is one option, but nonconvex constraints can hinder its convergence. As an alternative, we have developed a two-level distributed proportional fairness beamforming weight design (PFBWD) algorithm. This algorithm uses a combination of the augmented Lagrangian method (ALM) and a three-block ADMM framework. The proposed method effectively tackles nonconvexity, reduces complexity, and enables scalable, distributed optimization with guaranteed convergence.
Index Terms:
HAPS, vHetNet, interference management, cell-free, beamforming, proportional fairness, distributed algorithm
I Introduction
The next-generation (xG) wireless networks are evolving toward AI-driven, sensing-based systems that must deliver immersive throughput, hyper reliability, and low latency, while ensuring connectivity for everyone worldwide [1]. A wide range of innovative technologies is under consideration to meet these goals, spanning multiple layers of the telecommunication network, including terahertz frequency bands, movable antenna architectures, holographic beamforming, semantic communication, and advanced AI-driven resource management strategies. Meeting the ambitious performance and coverage demands requires more than just novel technologies. This also demands innovative network architectures, coupled with efficient performance optimization algorithms that can fully unlock their potential [2]. Among these architectural innovations, non-terrestrial networks (NTN), including space satellites, high altitude platform stations (HAPS), and uncrewed aerial vehicles, represent a particularly promising direction [3]. Unlike earlier deployments that primarily targeted rural and remote areas, future NTN will be tightly integrated with urban terrestrial networks to provide global coverage with seamless quality of experience.
In future xG wireless networks, HAPS is expected to play a critical role, particularly in urban environments with high user density and traffic demand. Positioned quasi-stationary in the stratosphere layer of the Earth at an altitude of approximately 20 km above the ground, HAPS offers several advantages over both space-based satellites and terrestrial networks. For instance, HAPS provides lower latency compared to space satellites, greater flexibility to support advanced technologies (e.g., ultra massive multiple-input multiple-output (umMIMO), reconfigurable intelligent surfaces), and a wider coverage area than terrestrial macro base stations (MBSs). As such, HAPS can serve as a complementary platform with the potential to help meet the demanding requirements of future networks, in conjunction with legacy terrestrial and satellite systems [4]. Moreover, the large surface area of HAPS enables the integration of ambient energy harvesting resources, contributing to the development of a more sustainable network architecture [5, 6]. Referred to by the International Telecommunication Union (ITU) as HAPS as International Mobile Telecommunications (IMT) base stations, HAPS can be integrated with existing terrestrial networks in urban environments to enhance coverage and capacity, thereby supporting the fulfillment of future network KPIs. Such integrated HAPS-terrestrial networks are referred to as HAPS-empowered vertical heterogeneous networks (vHetNets) [7].
Various tiers in vHetNets can operate over the same or different frequency bands [8]. Notably, the World Radiocommunication Conference 2023 (WRC-23) allocated additional IMT-identified spectrum to HAPS [9]. This development is pivotal for enabling the integration of HAPS with terrestrial wireless networks under a harmonized spectrum framework, where both HAPS and terrestrial tiers share the same frequency band. While such spectrum sharing enhances spectral efficiency (SE), it also leads to performance degradation due to inter- and intra-tier interference propagation [7]. This challenge is further amplified by the wide coverage area of HAPS. The high altitude of HAPS enables vHetNets to provide wide-area coverage that includes multiple terrestrial MBSs. Consequently, effective coordination between HAPS and these MBSs is essential for delivering optimal service to user equipment (UEs). Unlike terrestrial MBSs, which primarily coordinate with their immediate neighbors, HAPS must interact with all MBSs within its coverage area. This requirement results in large-scale optimization problems with greater complexity, often making rapid convergence difficult to achieve. Moreover, efficient decision-making depends on exchanging network status information between MBS and HAPS, which introduces significant signaling overhead on communication links.
I-A Related works
Since the wireless channel is interference-limited, designing efficient interference management strategies has a long research history [10]. Recent studies have explored spectrum sharing and interference suppression techniques in integrated HAPS-terrestrial networks. To this end, the authors in [11] proposed an interference canceler and coordination mechanism to facilitate spectrum sharing between HAPS and terrestrial systems. In [12], the authors introduced a cell design method for HAPS aimed at extending coverage while maintaining coexistence with terrestrial mobile networks. In addition, [13] presented a codebook-based interference suppression approach for space-air-ground integrated networks, demonstrating improved performance in managing inter-tier interference. In our previous work [4], we developed an interference management algorithm for an integrated HAPS-terrestrial network, leveraging subcarrier allocation and power control. Furthermore, in [14] and [15], we formulated three interference management problems aimed at optimizing user association and beamforming weights under different objective functions: weighted sum rate, max-min fairness, and proportional fairness (PF). The performance of the proposed algorithms was subsequently evaluated in both vHetNets and standalone terrestrial network scenarios. More recently, the authors in [16] proposed an interference mitigation framework to enable NTN and terrestrial network coexistence.
Among the existing approaches for managing interference, beamforming weight design has emerged as an efficient strategy [17]. To this end, optimal beamforming weights design algorithms for MIMO systems are developed in a centralized manner [18, 19]. However, due to the high computational and signaling overheads associated with centralized algorithms, significant efforts have also been devoted toward developing distributed algorithms. In  [20], the authors introduced an alternating direction method of multipliers (ADMM)-based beamforming approach for multi-group multicast systems, and in [21], the authors developed a stochastic ADMM framework for coordinated multi-cell beamforming. In addition, the authors in [22] proposed an ADMM-based distributed beamforming algorithm to minimize total transmit power under signal-to-interference-and-noise ratio (SINR) constraints in multi-cell networks. Moreover, [23] applied ADMM to power-efficient downlink beamforming in cell-free massive MIMO, also targeting SINR-constrained optimization across distributed access points.
Although the two-block ADMM approach has been widely adopted in distributed algorithm design [24], its limitations hinder its applicability to general nonconvex problems. To address this, more sophisticated methods have been explored. In [25], the authors formulated a joint user association and beamforming weight design problem aiming at maximizing the ergodic sum rate. Accordingly, the SINR term has been approximated with the average signal-to-leakage-and-noise ratio, and a three-step Gaussian belief propagation-based distributed solver is proposed to solve the problem. In [26], the authors formulated a sum rate maximization problem and introduced a belief propagation algorithm for cooperative MIMO systems that enables efficient downlink beamforming through decentralized message passing. In recent years, deep learning-based approaches have also gained traction. The authors in [27] employ deep reinforcement learning for distributed uplink beamforming in cell-free networks, considering a sum rate objective function. Similarly, [28] uses distributed graph-based learning to jointly optimize user association and beamforming in multi-cell networks to maximize the sum rate. Collectively, these contributions underscore the growing emphasis on distributed beamforming frameworks for xG wireless networks.
I-B Motivations and Contributions
It is evident that centralized algorithms favor small-scale systems; hence, they are practically infeasible in HAPS-empowered vHetNets, which are large-scale systems for several reasons. First, the large coverage area of HAPS encompasses a large number of MBSs, making it a large-scale system. Second, the use of ultra mMIMO involves a large number of antenna elements, resulting in a high-dimensional optimization problem. Therefore, the development of a distributed algorithm for beamforming design becomes necessary for handling large-scale networks such as HAPS-empowered vHetNets. Further, the formulated problems in vHetNets are inherently nonconvex, especially in interference-dominant networks [7]. To address this, various reformulation linearization techniques (RLT) are often employed to approximate the original problem and solve it iteratively through successive convex approximation (SCA) [29]. However, such iterative methods require solving multiple optimization problems iteratively, making them challenging for real-time or low-latency applications in dynamic networks. This reinforces the necessity of developing distributed algorithms tailored for nonconvex interference management problems in vHetNets. That said, directly applying conventional distributed approaches such as two-block ADMM to general nonconvex constrained problems does not guarantee convergence unless relaxation schemes are used [30]. To this end, multiple variants of ADMM approaches have been developed to address the nonconvexity of the optimization problems [31]. In this regard, the authors in [32] proposed a two-level distributed algorithm that can solve the nonconvex constrained problem in a distributed manner, with guaranteed convergence. Specifically, they introduce a key reformulation that guarantees the conditions for convergence of multi-block ADMM algorithms. The resulting two-level algorithm embeds a specially structured three-block ADMM in the inner loop within an augmented Lagrangian method (ALM) framework [32]. This algorithm is selected for its demonstrated advantages in accelerating convergence and improving feasibility in challenging nonconvex settings. Its ability to handle multi-block constraints robustly makes it well-suited for the problem structure.
In this work, we consider a HAPS-empowered vHetNet in which the base stations (BSs) [^1] The term BS in this work collectively refers to HAPS and MBSs. are equipped with MIMO antenna arrays. In [15], we demonstrated the performance gains achieved by associating UEs with multiple BSs. Moreover, the results in [15] showed that the PF objective function offers a fair trade-off between maximizing the sum rate and improving the worst-case UE SE. Therefore, we formulate a cell-free proportional fairness beamforming weight design (PFBWD) problem. The beamforming weights are designed to determine the strength of the signal received by each UE from each BS. Due to the problem’s nonconvex nature, directly applying the ADMM-based optimization approaches doesn’t guarantee convergence. To address this, we reformulate the original problem into an equivalent distributed form and then apply the two-level distributed algorithm. Accordingly, the main contributions of this paper are summarized as follows:
  • We adopt a cell-free approach as the underlying network architecture for HAPS-empowered vHetNet in an urban setting, where the HAPS coverage overlaps with the terrestrial network. A PFBWD problem is formulated to optimize network-wide fairness.
  • We apply the RLT to address the nonconvex nature of the PFBWD problem, and restructure the original problem into an equivalent distributed form.
  • We develop a novel two-level distributed algorithm to solve the large-scale PFBWD problem. The proposed algorithm incorporates a structured three-block ADMM at the inner-level within an outer-level ALM framework.
  • We theoretically establish the conditions required for the convergence of the proposed algorithm.
  • We validate and demonstrate the effectiveness of the proposed two-level distributed algorithm through numerical simulations, comparing it to its centralized counterpart in terms of SE and computational complexity.
[…]
The remainder of the paper is organized as follows. Section II introduces the system model and formulates the PFBWD optimization problem. Section III provides a distributed reformulation of the original nonconvex problem. Section IV describes the developed approach and algorithm. Section V discusses the convergence properties and computational complexity of the algorithm. Section VI presents and analyzes the numerical results, and finally, Section VII concludes the paper.
II System Model and Problem Formulation
We consider a cell-free HAPS-empowered vHetNet consisting of one HAPS and $B$ MBSs, simultaneously serving $U$ single antenna UEs as illustrated in Fig. 1. The vHetNet operates in the downlink channel within the sub-6 GHz frequency band. UEs and BSs are indexed by $u\in\mathcal{U}\triangleq\{1,\ldots,U\}$, and $b\in\mathcal{B}\triangleq\{1,\ldots,B+1\}$, respectively, where index $b=B+1$ is reserved for the HAPS. The MBSs and HAPS are equipped with MIMO antennas with a total number of $N_{b}=N^{H}_{b}\times N^{V}_{b}$ antenna elements in a uniform planar array configuration, where $N^{H}_{b}$ and $N^{V}_{b}$ are the number of antenna elements in the horizontal and vertical axes, respectively. We define $\mathbf{H}^{b}=[\mathbf{h}^{b}_{1},\dots,\mathbf{h}^{b}_{U}]\in\mathbb{C}^{N_{b}\times U}$, as the channel coefficients matrix at BS b, where $\mathbf{h}^{b}_{u}=[h^{b}_{1,u},\dots,h^{b}_{N_{b},u}]^{\top}\in\mathbb{C}^{N_{b}}$ denotes the channel vector between BS $b$ and UE $u$. We assume that all UEs share the same time-frequency resources with equal bandwidth allocation. Consequently, interference is propagated from all serving BSs to all UEs. To efficiently manage interference, BSs communicate with UEs using narrow and high-gain beams. Therefore, designing optimal beams is essential and the primary focus of this work.
Due to the strategic positioning of HAPS at an altitude of $20$ km above the ground, UEs experience a different channel compared to that of MBSs. Specifically, the propagation characteristics differ between the HAPS and terrestrial tiers, primarily because of the strong line of sight (LoS) link between HAPS and UEs. As a result, we will consider two distinct channel models, which are explained in detail below.
Figure 1: HAPS-empowered vHetNet system model.
II-A MBS-UE Channel
We consider small-scale fading, free-space path loss (FSPL), and shadowing for the channel between each UE and MBS. Particularly, the channel coefficient between the antenna element $r\in\{1,\dots,N_{b}\}$ of MBS $b$ and UE $u$, denoted as $h^{b}_{r,u}$, can be formulated as follows[33]:
$h^{b}_{r,u}=\frac{\overline{h}^{b}_{r,u}\xi^{b}_{u}}{\sqrt{{PL}_{b,u}}},~{}\forall r,~{}\forall u,~{}\forall b,$ (1)
where $\overline{h}_{r,u}$ represents the small-scale fading channel coefficient and follows Rayleigh distribution. $\xi^{b}_{u}=10^{{{\xi_{u}^{{}^{\prime}b}}}/10}$ denotes the log-normal shadowing gain, where ${\xi_{u}^{{}^{\prime}b}}$ is the Gaussian random variable with $\mathcal{N}(0,\sigma_{\xi}\,\text{dB})$. ${PL}_{b,u}$ refers to the FSPL between MBS $b$ and the UE $u$, and can me computed as
${{PL}_{b,u}}=\Bigg{(}\frac{4\pi f_{c}d_{b,u}}{c}\Bigg{)}^{2},~{}\forall b,~{}\forall u,$ (2)
where $f_{c}$ is the carrier frequency (in Hz), $d_{b,u}$ (in m) is the distance between the BS $b$ and the UE $u$, and $c$ represents the speed of light in free space.
II-B HAPS-UE Channel
Due to the high altitude of HAPS, there exist strong LoS links with a high probability to UEs. However, small-scale fading may still occur due to the multipath characteristics of urban environments. Particularly, $\mathbf{h}^{B+1}_{u}$, representing the channel coefficient vector between HAPS and UE $u$, can be modeled as a three-dimensional (3D) Rician fading channel with a dominant LoS and NLoS components as [34]
$\mathbf{h}_{u}^{B+1}=\frac{1}{\sqrt{{PL}_{(B+1),u}}}\Big{(}\sqrt{\frac{1}{1+K_{u}}}\overline{\mathbf{h}}_{u}+\sqrt{\frac{K_{u}}{1+K_{u}}}\overline{\mathbf{h}}_{u}\Big{)},~{}\forall u,\qquad\qquad\qquad$ (3)
where ${PL}_{(B+1),u}$ represents the FSPL between HAPS and UE $u$, calculated according to (2). $K_{u}$ is the Rician factor for UE $u$, and $\overline{\mathbf{h}}_{u}\in\mathbb{C}^{N_{B+1}}$ represents the NLoS component of the channel vector with its elements from a normal random distribution with zero mean and unit variance, $\mathcal{NC}(0,1)$. Accordingly, $\overline{\mathbf{h}}_{u}$ is the LoS component of the channel vector, given by [34] as
$\overline{\mathbf{h}}_{u}=\mathbf{a}(\theta_{u},\phi_{u})\otimes\mathbf{b}(\theta_{u},\phi_{u}),~{}\forall u,$ (4)
where
$\mathbf{a}(\theta_{u},\phi_{u})=[1,e^{j2\pi d_{h}},\dots,e^{j2\pi(N^{\text{H}}_{B+1}-1)d_{h}}]^{\top},~{}\forall u,$ (5a)
$\mathbf{b}(\theta_{u},\phi_{u})=[1,e^{j2\pi d_{v}},\dots,e^{j2\pi(N^{\text{v}}_{B+1}-1)d_{v}}]^{\top},~{}\forall u,$ (5b)
where $d_{h}=d_{\text{x}}\cos{\theta_{u}}\sin{\phi_{u}}/\lambda$ and $d_{v}=d_{\text{y}}\cos{\theta_{u}}\cos{\phi_{u}}/\lambda$. As depicted in Fig. 2, $\theta_{u}\in[0,\pi/2]$ and $\phi_{u}\in[-\pi,\pi)$ are the elevation and azimuth angles of UE $u$. Further, $d_{\text{x}}$ and $d_{\text{y}}$ are the antenna elements spacing in $y$ and $x$ directions and $\lambda=c/f_{c}$ represents the wavelength.
Figure 2: HAPS antenna architecture.
As mentioned earlier, we adopt the cell-free as the underlying network architecture where all UEs are served by all BSs. We assume that each BS has perfect information regarding the channel coefficients between antenna elements and UEs. Accordingly, the SINR of UE $u$ in the cell-free vHetNet can be expressed as [35]
$\gamma_{u}=\cfrac{\left|\sum_{b\in\mathcal{B}}\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{u}\right|^{2}}{\sum_{\mathcal{U}\setminus\{u\}}\left|\sum_{b\in\mathcal{B}}{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}}\right|^{2}+\sigma^{2}_{n}},~{}\forall u,$ (6)
where $\mathbf{w}^{b}_{u}\in\mathbb{C}^{N_{b}}$ refers to the beamforming vector for UE $u$ at BS $b$, and $\sigma^{2}_{n}$ is the variance of the additive white Gaussian noise (AWGN) at $u$. According to the above explanation, the preliminar PFBWD optimization problem can be expressed as
$\underset{\mathbf{W}^{b}}{\operatorname{maximize}}\,\,$ $~{}\sum_{u\in\mathcal{U}}{\log(\log_{2}(1+\gamma_{u}))}\,$
s.t. $\gamma_{u}\geq\gamma_{\text{min}},~{}\forall u,\qquad$
${\|\mathbf{W}^{b}\|^{2}_{F}}\leq P^{\text{max}}_{b},\,~{}\forall b,\qquad$
where $\mathbf{W}^{b}=[\mathbf{w}^{b}_{1},\dots,\mathbf{w}^{b}_{U}]\in\mathbb{C}^{N_{b}\times U}$ denotes the beamforming weight matrix at BS $b$, $P_{b}^{\text{max}}$ is the maximum allowable transmit power of BS $b$, and $\gamma_{\text{min}}$ is the minimum required SINR for each UE. In problem (II-B), constraint (II-B) ensures that each UE $u$ meets its minimum data rate requirement, while constraint (II-B) restricts the total transmit power of each BS $b$ to not exceed $P_{b}^{\text{max}}$. Problem (II-B) is nonconvex due to the objective function (II-B) and the nonconvex nature of constraint (II-B), rendering it intractable for conventional solvers. Typically, problems of this nature can be addressed by developing an iterative algorithm using the SCA framework. In each iteration of the SCA, a convex approximation of the original problem is solved. The solution obtained at convergence is considered suboptimal relative to the original nonconvex problem. Additionally, the approximated problems can be solved using either a centralized or a distributed algorithm. However, for large-scale problems, it is advisable to avoid a centralized algorithm, as it requires collecting channel coefficients from all BSs to solve the optimization problem, and thereafter distributing the decisions back to the BSs. This process imposes significant signaling overhead on the network. Additionally, the centralized approach must solve a large-scale optimization problem to determine the decision variables for the entire network, resulting in high computational complexity.
The problem at hand is large in scale and complexity, thereby necessitating the development of a distributed algorithm to solve it efficiently. To enable distributed computation, the original centralized formulation presented in (II-B) must first be transformed into an equivalent distributed formulation to which a distributed optimization method can be applied. The key objective of this reformulation is to enable parallel decomposition, whereby the problem is structured in a way that enforces per-BS constraints. This structure allows the overall problem to be divided and solved across multiple BSs simultaneously. In the subsequent section, we focus on this transformation process, aiming to recast the original problem (II-B) into a distributed format suitable for distributed implementation.
III Distributed Reformulation of the Original Problem (II-B)
In this section, we detail the series of reformulations required to transform the problem (II-B) to a distributed format that can be solved by the two-level distributed algorithm. To this end, first, we deal with the objective function (II-B). We define $U$ slack variables $t_{u},~{}\forall u,$ as the lower bound for $\log_{2}(1+\gamma_{u}),~{}\forall u$. Accordingly, the objective function can be replaced by $f(\mathbf{t})=\sum_{u=1}^{U}\log_{2}(t_{u})$, where $\mathbf{t}=[1,\dots,t_{U}]^{\top}\in{\mathbb{R}^{+}}^{U}$. The $U$ newly defined constraints will be added to the problem as
$t_{u}\leq\log_{2}(1+\gamma_{u}),~{}\forall u.$ (8)
Next, to deal with the nonconvexity involved in (II-B), we replace it equivalently as
$\alpha_{u}\beta_{u}\leq\left|\sum_{b\in\mathcal{B}}{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{u}}\right|^{2},~{}\forall u,$ (9a)
$\beta_{u}\geq\sum_{\mathcal{U}\setminus\{u\}}\left|\sum_{b\in\mathcal{B}}{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}}\right|^{2}+\sigma^{2}_{n},\,~{}\forall u,$
where $\alpha_{u},~{}\forall u,$ and $\beta_{u},~{}\forall u,$ are new slack variables representing the SINR $\gamma_{u}$, and the total interference-plus-noise of user $u$, respectively. Constraint (III) is convex, whereas constraint (9a) is nonconvex due to the presence of the norm function on its right-hand side and the product of two variables on the left-hand side. For now, we retain the nonconvex constraint (9a) as it is, since the two-level distributed algorithm is capable of handling it.
To facilitate the distributed restructuring of the problem, we revisit the right-hand side of constraint (III). Specifically, the current form of (III) begins with a summation over interfering UEs, whereas we prefer a formulation where the summation over BSs appears first to better align with the distributed structure. To this end, we apply Jensen’s inequality, which yields the following upper bound:
$\sum_{\mathcal{U}\setminus\{u\}}\left|\sum_{b\in\mathcal{B}}{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}}\right|^{2}\leq\sum_{b\in\mathcal{B}}\sum_{\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|^{2}},~{}\forall u.$ (10)
Accordingly, constraint (III) can be rewritten as
$\beta_{u}\geq\sum_{b\in\mathcal{B}}\sum_{\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|^{2}}+\sigma^{2}_{n},\,~{}\forall u.$ (11)
To solve the problem through the two-level distributed algorithm, we introduce a few more new slack variables $\mathbf{s}_{b}=[s_{b,1},\dots,s_{b,U}]^{\top}\in{\mathbb{R}}^{U}~{}\text{and}~{}\mathbf{I}_{b}=[I_{b,1},\dots,I_{b,U}]^{\top}\in{\mathbb{R}^{+}}^{U}$ to (13a) replace constraints (9a) and (11) equivalently as
$\alpha_{u}\beta_{u}\leq\left|\sum_{b\in\mathcal{B}}s_{b,u}\right|^{2},~{}\forall u,$ (12a)
$\beta_{u}\geq\sum_{b\in\mathcal{B}}{I_{b,u}}+\sigma^{2}_{n},\,~{}\forall u,$
where $s_{b,u}$ and $I_{b,u}$ are defined as follows:
$s_{b,u}\leq{\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{u}},~{}\forall u,~{}\forall b,$ (13a)
$I_{b,u}\geq\sum_{\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|^{2}},~{}\forall u,~{}\forall b.$
Furthermore, defining $\alpha_{u}$ as the lower bound for $\gamma_{u}$, constraints (II-B) and (8) will be replaced by (14a) and (III), respectively.
$\alpha_{u}\geq\gamma_{\text{min}},~{}\forall u,$ (14a)
$e^{t_{u}}\leq 1+\alpha_{u},~{}\forall u.$
Constraint (III) is convex; however, to reduce computational complexity, it can be reformulated as a series of second-order cone (SOC) constraints using an exponential cone approximation, as shown in (15a[36]. In this formulation, the parameter $\nu$ controls the accuracy of the approximation, and $k^{u}_{i},\ i=1,\dots,\nu+4,\ \forall u$, are newly introduced slack variables.
$k^{u}_{\nu+4}\leq 1+\alpha_{u},~{}\forall u,$ (15a)
$\|[2+{t_{u}/2^{(\nu-1)}}~{}~{}1-k^{u}_{1}]^{\top}\|_{2}\leq 1+k^{u}_{1},~{}\forall u,$
$\|[5/3+{t_{u}/2^{\nu}}~{}~{}1-k^{u}_{2}]^{\top}\|_{2}\leq 1+k^{u}_{2},~{}\forall u,$
$\|[2k^{u}_{1}~{}~{}1-k^{u}_{3}]^{\top}\|_{2}\leq 1+k^{u}_{3},~{}\forall u,$
$19/72+k^{u}_{2}+1/24k^{u}_{3}\leq k^{u}_{4},~{}\forall u,$
$\|[2k^{u}_{i-1}~{}~{}1-k^{u}_{i}]^{\top}\|_{2}\leq 1+k^{u}_{i},~{}\forall u,~{}i=5,\dots,\nu+4.\qquad$
After aforementioned transformations, problem (II-B) can be reformulated as
$\underset{\mathbf{W},\mathbf{t},\boldsymbol{\beta},\boldsymbol{\alpha},\mathbf{S},\mathbf{I}}{\operatorname{maximize}}\,\,$ $\sum_{u\in\mathcal{U}}{\log(t_{u})}\,$
[…]
where $\boldsymbol{\beta}$, $\boldsymbol{\alpha}$, $\mathbf{S}$, and $\mathbf{I}$ are vectors and matrices that collect the variables $\beta_{u}$, $\alpha_{u}$, $s_{b,u}$, and $I_{b,u},~{}\forall u,~{}\forall b$, respectively. As can be observed, the set of constraints (II-B) and (13a) are defined per BS, while the set of constraints (12a), (14a), and (15a) are defined per UE. However, these two sets of constraints are coupled through the variables $s_{b,u}$ and $I_{b,u}$. To address this coupling, we introduce new slack variables $\overline{\mathbf{s}}_{b}=[\overline{s}_{b,1},\dots,\overline{s}_{b,U}]^{\top}\in{\mathbb{R}^{+}}^{U}$ and $\overline{\mathbf{I}}_{b}=[\overline{I}_{b,1},\dots,\overline{I}_{b,U}]^{\top}\in{\mathbb{R}^{+}}^{U}$, which are associated with $\mathbf{s}_{b}$ and $\mathbf{I}_{b}$, respectively, such that they satisfy the following equality constraints:
$\mathbf{s}_{b}=\overline{\mathbf{s}}_{b},~{}\mathbf{I}_{b}=\overline{\mathbf{I}}_{b},~{}\forall b.$ (17)
Consequently, we replace $s_{b,u}$ and $I_{b,u}$ in the per BS constraints (13a) with $\overline{s}_{b,u}~{}\text{and}~{}\overline{I}_{b,u}$, as
$\overline{s}_{b,u}\leq\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{u},~{}\forall u,~{}\forall b,$ (18a)
$\overline{I}_{b,u}\geq\sum_{\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|}^{2},~{}\forall u,~{}\forall b.$
Now, problem (III) can be reformulated as
$\underset{\overline{\mathbf{S}},\overline{\mathbf{I}},\mathbf{W},\mathbf{t},\boldsymbol{\beta},\boldsymbol{\alpha},\mathbf{S},\mathbf{I}}{\operatorname{minimize}}\,\,$ $-\sum_{u\in\mathcal{U}}{\log(t_{u})}\,$
[…]
Problem (III) represents a distributed reformulation of the original problem (II-B). In this reformulation, the decision variables and constraints are decomposed into local components, associated with individual BSs, and global components that coordinate their interactions. This separation enables each BS to solve its local subproblem independently, subject to a set of local constraints, while only a limited amount of information must be exchanged to satisfy the global constraints. Such a structure facilitates parallel decomposition and allows the problem to be solved efficiently in a distributed manner. Accordingly, we define the compact variable sets as follows: $\mathcal{X}=\{\mathbf{t},\boldsymbol{\beta},\boldsymbol{\alpha},\mathbf{S},\mathbf{I}\}$, representing the set of global optimization variables, and $\overline{\mathcal{X}}=\{\overline{\mathbf{S}},\overline{\mathbf{I}},\mathbf{W}\}$, denoting the set of local optimization variables. Accordingly, the associated constraint sets are defined as
[…]
where $\mathcal{C}$ represents a nonconvex constraint set, and $\overline{\mathcal{C}}$ is a compact convex set. These sets are coupled through the equality constraints defined in (17). The problems in the form of (III) can be expressed in the general distributed form as
$\underset{\mathcal{X},\overline{\mathcal{X}}}{\operatorname{minimize}}\,\,$ $f(\mathcal{X})+g(\overline{\mathcal{X}})\,$
s.t. $\mathbf{A}\mathcal{X}+\mathbf{B}\overline{\mathcal{X}}=0,$
$\mathcal{X}\in\mathcal{C}.$
It is worth noting that some common distributed algorithms, such as the two-block ADMM, can be employed to solve problems in the form of (III). However, the convergence of the two-block ADMM algorithm relies on the following two conditions:
  • Condition 1: $\operatorname{Im}(\mathbf{A})\subseteq\operatorname{Im}(\mathbf{B})$, where $\operatorname{Im}(.)$ denotes the image space of the argument matrix. [^2] The image space (or column space) of a matrix $\mathbf{A}$ is the set of all vectors that can be expressed as $\mathbf{A}\mathcal{X}$ for some vector $\mathcal{X}$.
  • Condition 2: The last-block optimization variables are unconstrained and the objective function $g(\overline{\mathcal{X}})$ is Lipschitz differentiable. [^3] A function $f$ is Lipschitz differentiable if it is differentiable and its gradient is Lipschitz continuous, meaning the gradient does not change too rapidly.
In problem (III), the Condition 2 does not hold. Specifically, the last-block variable $\overline{\mathcal{X}}$ cannot be structured in a way that ensures Lipschitz differentiability. This difficulty arises because the variable set $\overline{\mathcal{X}}$ is constrained by a nonconvex set $\overline{\mathcal{C}}$. Even if the constraints defined by $\overline{\mathcal{C}}$ were relaxed and incorporated into the objective function through penalty terms, the resulting formulation would still yield a nonconvex and nonsmooth objective. Consequently, the two-block ADMM approach does not guarantee convergence when applied to problem (III). To overcome this limitation, the next section introduces the proposed distributed PFBWD algorithm, which adopts the two-level distributed framework developed in [32] to efficiently solve problem (III). This framework integrates a structured three-block ADMM method within an ALM-based outer loop, enabling scalable and efficient distributed optimization. Further details of the proposed algorithm are provided in the following section.
IV Proposed Two-Level Distributed PFBWD Algorithm
To develop the two-level distributed algorithm, we first reformulate problem (III) to make it suitable for the application of the three-block ADMM approach. Then, we design a two-level iterative algorithm, consisting of inner- and outer-level, to solve the reformulated problem. The inner-level algorithm utilizes the three-block ADMM method, with its iterations represented by $t$. Whereas, the outer-level algorithm employs the ALM, with its iterations represented by $k$, to update the Lagrange multipliers and enforce the equality constraints. The aforementioned procedure is described in detail in the following subsections.
IV-A Reformulation of Problem (III)
For reformulation, we introduce auxiliary variables $\mathbf{z}_{l,b}\in\mathbb{R}^{U},~{}l\in\mathcal{L}\triangleq\{s,I\},~{}\forall b,$ and replace the equality constraints in (17) with the following constraints as
$\mathbf{s}_{b}-\overline{\mathbf{s}}_{b}+\mathbf{z}_{s,b}=0,\quad\forall b,$ (22a)
$\mathbf{I}_{b}-\overline{\mathbf{I}}_{b}+\mathbf{z}_{I,b}=0,\quad\forall b,$
To ensure equivalence with the original problem (III), we additionally enforce the following constraints:
$\mathbf{z}_{l,b}=0,~{}\forall l,~{}\forall b.$ (23)
The introduction of auxiliary variables, $\mathbf{z}_{l,b}$, yields two important structural properties:
  • Property I: The reformulated linear coupling constraints (22a) now involve three blocks of variables. The third block corresponds to auxiliary variables associated with an identity matrix, whose image spans the full space. This structure satisfies Condition 1.
  • Property II: If the constraints in (23) are temporarily ignored, the problem becomes compatible with the standard three-block ADMM framework. In this case, the third block variable is unconstrained, thereby satisfying Condition 2 required for ADMM convergence.
To facilitate Property 2, we relax constraint (23) by using ALM method, and thus, the objective function can now be written as
$f(\mathcal{X},\mathcal{Z})\!=\!-\sum_{u\in\mathcal{U}}\log(t_{u})+\sum_{b\in\mathcal{B}}\!\sum_{l\in\mathcal{L}}\!\left(\boldsymbol{\lambda}^{\top}_{l,b}\mathbf{z}_{l,b}+\frac{\rho_{o}}{2}\|\mathbf{z}_{l,b}\|^{2}\right),$ (24)
where $\mathcal{Z}$ denotes collection of auxiliary variables $\mathbf{z}_{l,b}$, $\boldsymbol{\lambda}_{l,b}$ represents the Lagrange multipliers associated with the constraints $\mathbf{z}_{l,b}=0$, and $\rho_{o}>0$ is the penalty parameter for the outer-level ALM. Therefore, problem (III) is reformulated as
$\underset{\mathcal{X},\overline{\mathcal{X}},\mathcal{Z}}{\operatorname{minimize}}$ $~{}f(\mathcal{X},\mathcal{Z})$
[…]
The augmented Lagrangian term in (24) is both strongly convex in $\mathcal{Z}$ and Lipschitz differentiable, which simultaneously satisfies Conditions 1 and 2, and hence, ensures the convergence of the three-block ADMM algorithm to solve problem (IV-A). However, the solution to the above problem does not necessarily ensure that constraint (23) would be satisfied. To this end, we rely on ALM method which ensures variables $z_{l,b}$ to go to zero by updating $\lambda_{l,b}$. Specifically, this reformulation decomposes the original problem (III) into a two-level structure. The outer-level ALM algorithm iteratively drives the slack variables $\mathcal{Z}$ toward zero by updating the multipliers $\boldsymbol{\lambda}_{l,b}$ at each outer iteration indexed by $k$, thereby ensuring convergence to a stationary point of the original problem (III[32]. Whereas, the inner-level algorithm solves the reformulated problem using the three-block distributed ADMM while ensuring convergence guarantees. The details of the inner-level algorithm are presented in the following subsection.
IV-B Inner-Level Three-Block ADMM Algorithm
Given $\boldsymbol{\lambda}^{(k)}_{l,b}$ and $\rho^{(k)}_{o}$ as the Lagrange multipliers and penalty parameter for the $k$th outer-level iteration, the augmented Lagrangian function corresponding to the $k$th inner-level problem is defined in (26).
$L(\mathcal{X},\overline{\mathcal{X}},\mathcal{Z},\boldsymbol{\Psi})=f(\mathcal{X},\mathcal{Z})+\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\left({\boldsymbol{\psi}^{\top}_{l,b}}\mathbf{r}_{l,b}+\frac{\rho_{l,b}}{2}\|\mathbf{r}_{l,b}\|^{2}\right)=-\sum_{u\in\mathcal{U}}{\log(t_{u})}+\sum_{b\in\mathcal{B}}{L^{b}(\mathcal{X}_{b},\overline{\mathcal{X}}_{b},\mathbf{Z}_{b},\boldsymbol{\psi}_{b})},$ (26)
where:
$L^{b}(\mathcal{X}_{b},\overline{\mathcal{X}}_{b},\mathbf{Z}_{b},\boldsymbol{\psi}_{b})$$\triangleq$ $\sum_{l\in\mathcal{L}}\left(\boldsymbol{\lambda}^{\top}_{l,b}\mathbf{z}_{l,b}+\frac{\rho_{o}}{2}\|\mathbf{z}_{l,b}\|^{2}+{\boldsymbol{\psi}^{\top}_{l,b}}\mathbf{r}_{l,b}+\frac{\rho_{l,b}}{2}\|\mathbf{r}_{l,b}\|^{2}\right),~{}\forall b.$ (27)
In (26) and (27), matrices $\mathbf{Z}_{b},~{}\forall b,$ include the vectors $\mathbf{z}_{l,b},~{}\forall l$, and $\boldsymbol{\psi}_{l,b}$ denotes the Lagrange multipliers associated with the equality constraints in (22a), which are reflected by the defined residuals $\mathbf{r}_{l,b}$, as
$\mathbf{r}_{s,b}=\mathbf{s}_{b}-\overline{\mathbf{s}}_{b}+\mathbf{z}_{s,b},~{}\forall b,$ (28a)
$\mathbf{r}_{I,b}=\mathbf{I}_{b}-\overline{\mathbf{I}}_{b}+\mathbf{z}_{I,b},~{}\forall b.$
According to the three-block ADMM approach, problem (IV-A) can be partitioned into three variable blocks and solved iteratively in an alternative manner until the convergence criteria are met. The three blocks, and the Lagrange multipliers are updated according to the following equations:
ADMM Block 1: (29a)
$\quad\underset{\mathcal{X}\in\mathcal{C}}{\operatorname{minimize}}\,~{}L(\mathcal{X},\overline{\mathcal{X}}^{(t-1)},\mathcal{Z}^{(t-1)},\boldsymbol{\Psi}^{(t-1)})$ (29b)
ADMM Block 2:
$\quad\underset{\overline{\mathcal{X}}\in\overline{\mathcal{C}}}{\operatorname{minimize}}\,~{}L(\mathcal{X}^{(t)},\overline{\mathcal{X}},\mathcal{Z}^{(t-1)},\boldsymbol{\Psi}^{(t-1)})$ (29c)
ADMM Block 3:
$\quad\underset{\mathcal{Z}}{\operatorname{minimize}}\,~{}L(\mathcal{X}^{(t)},\overline{\mathcal{X}}^{(t)},\mathcal{Z},\boldsymbol{\Psi}^{(t-1)})$ (29d)
Lagrange Multiplier Update:
$\quad\boldsymbol{\psi}^{(t)}_{l,b}\leftarrow\boldsymbol{\psi}^{(t-1)}_{l,b}+\rho^{(k)}_{l,b}\mathbf{r}^{(t)}_{l,b},~{}\forall l,~{}\forall b.$ (29e)
In the remainder of this section, we provide a detailed explanation of the optimization problem for each block, followed by the pseudo code of the entire inner-level procedure in Algorithm 2.
IV-B1 Inner-level ADMM block 1 optimization problem
The optimization problem solved in the $t$-th iteration of the inner-level algorithm block 1 is given by (29b). In problem (29b), the constraint set $\mathcal{C}$ includes the nonconvex constraint (12a). Consequently, problem (29b) cannot be solved directly by standard solvers and requires either convex approximations or iterative resolution via SCA. However, since the problem size is small, the iterative algorithm is very fast. To facilitate this, we introduce new slack variables $\mathbf{p}=[p_{1},\dots,p_{U}]^{\top}\in\mathbb{R}^{U}$ and $\mathbf{q}=[q_{1},\dots,q_{U}]^{\top}\in\mathbb{R}^{U}$, representing the real and imaginary parts, respectively, of the summation on the right-hand side of (12a), defined as
$p_{u}=\Re{\left(\sum_{b\in\mathcal{B}}s_{b,u}\right)},~{}q_{u}=\Im{\left(\sum_{b\in\mathcal{B}}s_{b,u}\right)},~{}\forall u.$ (30)
In this way, the right hand side of inequality (12a) will be replaced with by $p_{u}^{2}+q_{u}^{2}$. Accordingly, moving $\beta_{u}$ to the right hand side and replacing the right hand side with its approximated first-order Taylor series, the constraint can be approximated as
${}\alpha_{u}\leq\frac{2{p^{(m)}_{u}}}{{\beta^{(m)}_{u}}}\left(p_{u}-{p^{(m)}_{u}}\right)+\frac{2{q^{(m)}_{u}}}{{\beta^{(m)}_{u}}}\left(q_{u}-{q^{(m)}_{u}}\right)+\frac{\left({p^{(m)}_{u}}\right)^{2}+\left({q^{(m)}_{u}}\right)^{2}}{\beta^{(m)}_{u}}\left(1-\frac{\beta_{u}-{\beta^{(m)}_{u}}}{{\beta^{(m)}_{u}}}\right),~{}\forall u,$ (31)
where the superscript $(m)$ refers to the value of the corresponding variable in the $m$-th iteration of the SCA process. In this way, the $m$-th SCA iteration equivalent problem for problem (29b) will be as
$\underset{\mathcal{X}}{\operatorname{minimize}}\,\,$ $L(\mathcal{X},\overline{\mathcal{X}}^{(t-1)},\mathcal{Z}^{(t-1)},\boldsymbol{\Psi}^{(t-1)})\,$
[…]
Algorithm 1 describes the first-block SCA-based algorithm. Algorithm 1 solves the approximated problem (IV-B1) iteratively for the optimal solution, which is the suboptimal solution of the original problem (29b). The algorithm stops upon convergence, which is defined as when a tolerable change in the objective function value is achieved (less than $\epsilon$), or the maximum number of SCA iterations, $N_{\text{iter}}$, is reached, whichever first. This algorithm can be solved centrally or distributedly at each BS using the same initial and broadcasted values.
Algorithm 1 SCA-based algorithm for ADMM Block 1.
1:  Initialize $\mathbf{p}^{(0)}$, $\mathbf{q}^{(0)}$, and $\boldsymbol{\beta}^{(0)},$ set $m\leftarrow 0$.
2:  repeat
3:     Solve (IV-B1) to find $\mathcal{X}^{(m)*}$, $\mathbf{p}^{(m)*}$, $\mathbf{q}^{(m)*}$.
4:     Update $\mathbf{p}^{(m+1)}\leftarrow\mathbf{p}^{(m)*}$, $\mathbf{q}^{(m+1)}\leftarrow\mathbf{q}^{(m)*}$, $m\leftarrow m+1$.
5:  until convergence or $m>N_{\text{iter}}$.
IV-B2 Inner-level ADMM block 2 optimization problem
The optimization problem to be solved in ADMM Block 2 is given in (29c). It is a convex problem that can be solved using standard solvers. Moreover, the structure of ADMM Block 2 lends itself well to decomposition into $(B+1)$ subproblems, each of which can be solved locally by a BS in parallel. In particular, the constraint set $\overline{\mathcal{C}}$ imposes per BS constraints, and the objective function $L(\mathcal{X}^{(t)},\overline{\mathcal{X}},\mathcal{Z}^{(t-1)},\boldsymbol{\Psi}^{(t-1)})$ is the sum of independent per–BS terms. Therefore, we can separate the objective and constraints so that each BS solves its subproblem in parallel. In this way, the optimization problem at each BS $b$ becomes
$\underset{\overline{\mathcal{X}}_{b}}{\operatorname{minimize}}\,\,$ $L^{b}(\mathcal{X}^{(t)}_{b},\overline{\mathcal{X}}_{b},\mathbf{Z}^{(t-1)}_{b},\boldsymbol{\psi}^{(t-1)}_{b})\,$
s.t. ${\|\mathbf{W}^{b}\|^{2}_{F}}\leq P^{\text{max}}_{b},$
$\overline{s}_{b,u}\leq\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{u},~{}\forall u,$
$\overline{I}_{b,u}\geq\sum_{\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|}^{2},~{}\forall u,$
where $L^{b}(\mathcal{X}^{(t)}_{b},\overline{\mathcal{X}}_{b},\mathbf{Z}^{(t-1)}_{b},\boldsymbol{\psi}^{(t-1)}_{b})$ is defined according to (27).
IV-B3 Inner-level ADMM block 3 optimization problem
The third block of the Inner-level ADMM algorithm, at iteration $t$, solves the problem (29d) to obtain the $\mathcal{Z}^{(t)}$ values. Problem (29d) represents a convex optimization problem, and with the same logic as block 2, can be decomposed to $(B+1)$ independent optimization problems which can be solved in parallel by each BS $b$. The ADMM block 3 optimization problem to be solved at each BS $b$ will be as
$\underset{\mathbf{Z}_{b}}{\operatorname{minimize}}\,\,$ $L^{b}(\mathcal{X}^{(t)}_{b},\overline{\mathcal{X}}^{(t)}_{b},\mathbf{Z}_{b},\boldsymbol{\Psi}^{(t-1)}_{b})\,$ (34a)
Algorithm 2 The $k$th inner-level ADMM Algorithm.
1:  Input: $\mathbf{I}^{(k-1)}$, $\mathbf{S}^{(k-1)}$, $\mathcal{Z}^{(k-1)}$, $\boldsymbol{\lambda}_{l,b}^{(k)},~{}\rho_{o}^{(k)}$.
2:  Output: $\{\mathbf{W}^{b*}\}_{b=1}^{B+1}$, $\mathbf{I}^{(k)}$, $\mathbf{S}^{(k)}$, and $\mathcal{Z}^{(k)}$.
3:  Initialization:
4:         $t\leftarrow 1$; $n\leftarrow 1$, $\rho^{(k)}_{l,b}=\delta\rho^{(k)}_{{o}}$.
5:         $\mathbf{I}^{(0)}\leftarrow\mathbf{I}^{(k-1)}$, $\mathbf{S}^{(0)}\leftarrow\mathbf{S}^{(k-1)}$, $\mathcal{Z}^{(0)}\leftarrow\mathcal{Z}^{(k-1)}$.
6:         Choose $\boldsymbol{\psi}_{l,b}^{(0)},~{}\forall b,$ such that:
7:               $\boldsymbol{\lambda}_{l,b}^{(k)}+\rho_{o}^{(k)}\mathbf{z}_{l,b}^{(0)}+\boldsymbol{\psi}_{l,b}^{(0)}=0$ for $l\in\mathcal{L},~{}\forall b$.
8:         Set tolerance thresholds $\{\epsilon_{1},\epsilon_{2},\epsilon_{l,b}\}$.
9:  while stopping criteria (35a) is not met do
10:     for all BSs $b\in\mathcal{B}$ (in parallel) do
11:        ADMM Block 1:    Solve the problem (IV-B1) using Algorithm 1.
12:        ADMM Block 2:    Solve problem (IV-B2).
13:        ADMM Block 3:    Solve problem (34a).
14:        Lagrange Multiplier Update:
15:           $\boldsymbol{\psi}^{(t)}_{l,b}\leftarrow\boldsymbol{\psi}^{(t-1)}_{l,b}+\rho^{(k)}_{l,b}\mathbf{r}^{(t)}_{l,b},~{}\forall l.$
16:        Broadcast $\mathbf{s}_{b}^{(t)}$, $\mathbf{I}_{b}^{(t)}$, and $\mathbf{Z}_{b}^{(t)},$ to all BSs.
17:     end for
18:     $t\leftarrow t+1$
19:  end while
20:  Return $\mathbf{I}^{(t)}$, $\mathbf{S}^{(t)}$, $\mathcal{Z}^{(t)}$.
Algorithm 2 summarizes the $k$-th inner-level three-block ADMM procedure for solving the distributed PFBWD problem. At each outer-level iteration indexed by $k$, the algorithm initializes local variables and Lagrange multipliers, then alternatively updates the three variable blocks, $\mathcal{X}$, $\overline{\mathcal{X}}$, and $\mathcal{Z}$, while enforcing the equality constraints via augmented Lagrangian terms. This structure enables the BSs to solve their local subproblems in parallel during Blocks 2 and 3. This is further explained in detail later in this section. Accordingly, the Lagrange multipliers are updated using the residuals of the coupling constraints, and few information is exchanged among the BSs. The procedure repeats until the specified stopping criteria are satisfied, at which point the updated variables are returned for use in the next outer-level ALM iteration. According to [32], Algorithm 2 terminates once it identifies a point $(\mathcal{X}^{(t)},\overline{\mathcal{X}}^{(t)},\mathcal{Z}^{(t)})$ that satisfies the stopping conditions specified in (35a). The rationale behind the stopping criteria (35a) will be discussed in Section V.
$\left\|\sum_{b\in\mathcal{B}}\Bigl{[}\rho^{(k)}_{s,b}\bigl{(}-\overline{\mathbf{s}}^{(t)}_{b}+\mathbf{z}^{(t)}_{s,b}+\overline{\mathbf{s}}^{(t-1)}_{b}-\mathbf{z}^{(t-1)}_{s,b}\bigr{)}+\rho^{(k)}_{I,b}\bigl{(}-\overline{\mathbf{I}}^{(t)}_{b}+\mathbf{z}^{(t)}_{I,b}+\overline{\mathbf{I}}^{(t-1)}_{b}-\mathbf{z}^{(t-1)}_{I,b}\bigr{)}\Bigr{]}\right\|\leq\epsilon_{1},\qquad$ (35a)
$\left\|\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\Bigl{[}\rho^{(k)}_{l,b}\bigl{(}\mathbf{z}^{(t-1)}_{l,b}-\mathbf{z}^{(t)}_{l,b}\bigr{)}\Bigr{]}\right\|\leq\epsilon_{2},\qquad$
$\left\|\mathbf{r}^{(t)}_{l,b}\right\|\leq\epsilon_{l,b},~{}\forall l,~{}\forall b.$
IV-C Proposed Distributed Two-Level PFBWD Algorithm
In this section, we integrate the previously described components to present the proposed distributed two-level PFBWD algorithm. As shown in Algorithm 3, the algorithm is designed for a HAPS-empowered vHetNet consisting of $B$ MBSs and one HAPS. It adopts an outer–inner structure, where the outer-level updates the Lagrange multipliers and penalty parameters, while the inner-level, executed via the three-block ADMM in Algorithm 2, handles the main optimization. At each outer iteration $k$, the central controller broadcasts initial variables to all BSs. Each BS then solves its local subproblem in parallel to update local variables and interference estimates. The central controller collects these updates to adjust the Lagrange multipliers and the penalty parameter based on the residual norm, thereby enhancing convergence. The proposed approach is both efficient and scalable, as it exploits problem separability across BSs to enable parallel computation without requiring channel state information exchange with the central node. This significantly reduces signaling overhead, which is especially beneficial in networks with large-scale antenna arrays, particularly at the HAPS. The stopping criteria for both inner- and outer-levels, along with a complexity analysis, are provided in Section V.
Algorithm 3 Proposed Two-Level Distributed PFBWD Algorithm for a HAPS-Empowered vHetNet with $B$ MBSs and One HAPS.
1:  Input: $U$$B$$\sigma^{2}_{n}$.
2:  Output: $\{\mathbf{W}^{b*}\}_{b=1}^{B+1}$
3:  Initialize:
4:      $\mathbf{S}^{(0)},~{}\mathcal{Z}^{(0)},~{}\boldsymbol{\lambda}_{l,b}^{(0)}\in\mathbb{R}^{(B+1)\times U},~{}\forall l,~{}\forall b$,
5:      $\mathbf{I}^{(0)}\in\mathbb{R}_{+}^{(B+1)\times U}$$\rho^{(1)}_{o}>0,~{}\omega\in[0,1),~{}\gamma>1,~{}k\leftarrow 1$
6:  while stopping criterion is not met do
7:     Broadcast current variables to all BSs.
8:     Use $(\mathbf{S}^{(k-1)},\mathbf{I}^{(k-1)},\mathcal{Z}^{(k-1)})$ as the initial point and solve Algorithm 2.
9:     Obtain updated values $(\mathbf{S}^{(k)},\mathbf{I}^{(k)},\mathcal{Z}^{(k)})$.
10:     for all BSs $b\in\mathcal{B}$ (in parallel) do
11:        for all $l\in\mathcal{L}$ do
12:           $\boldsymbol{\lambda}_{l,b}^{(k+1)}\leftarrow\text{Proj}\bigl{(}\boldsymbol{\lambda}_{l,b}^{(k)}+\rho^{(k)}_{o}\mathbf{z}_{l,b}^{(k)}\bigr{)}$
13:        end for
14:     end for
15:     if $\|\mathbf{Z}_{b}^{(k)}\|\geq\omega\|\mathbf{Z}_{b}^{(k-1)}\|,~{}\forall b,$ then
16:        $\rho^{(k+1)}_{o}\leftarrow\gamma\rho^{(k)}_{o}$
17:     end if
18:     $k\leftarrow k+1$
19:  end while
V Convergence and Computational Complexity Analysis
In this section, we will provide details regarding the convergence and computational complexity analysis of the proposed distributed PFBWD algorithm.
V-A Stopping Criteria
V-A1 Stopping Criteria for the inner-level algorithm 2
The $k$-th inner-level algorithm aims to find an approximate stationary solution $(\mathcal{X}^{(k)},\overline{\mathcal{X}}^{(k)},\mathcal{Z}^{(k)})$ of the problem (IV-A). Therefore, defining $f(\mathcal{X})=-\sum_{u\in\mathcal{U}}\log(t_{u})$, according to the first order optimality condition of constrained optimization problems [37], at convergence of the $t$-th inner-level ADMM problem, the generated point $(\mathcal{X}^{(k)},\overline{\mathcal{X}}^{(k)},\mathcal{Z}^{(k)})$ should satisfy the first order optimality condition of the problem (IV-A) in the sense that there exist $d^{(k)}_{j},~{}\forall j\in\{1,2\},~{}d^{(k)}_{3,b},~{}\text{and}~{}d^{(k)}_{l,b}$, such that
$d^{(k)}_{1}\in\nabla f(\mathcal{X}^{(k)})+\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(k)}_{l,b}+N_{\mathcal{C}}(\mathcal{X}^{(k)}),\qquad$ (36a)
$d^{(k)}_{2}\in-\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(k)}_{l,b}+N_{\overline{\mathcal{C}}}(\overline{\mathcal{X}}^{(k)}),\qquad$
$0=\mathbf{\lambda}^{(k)}_{l,b}+\rho^{(k)}_{{o}}\mathbf{z}^{(k)}_{l,b}+\boldsymbol{\psi}^{(k)}_{l,b},~{}\forall b,~{}\forall l,\qquad$
$d^{(k)}_{l,b}=\mathbf{r}^{(k)}_{l,b},~{}\forall l,~{}\forall b,\qquad$
$\|d^{(k)}_{j}\|\leq\epsilon^{(k)}_{j},~{}\forall j\in\{1,2\},~{}\|d^{(k)}_{l,b}\|\leq\epsilon^{(k)}_{l,b},~{}\forall l,~{}\forall b,\qquad$
where $\epsilon^{(k)}_{j},\epsilon^{(k)}_{l,b},~{}\forall j\in\{1,2\},~{}\forall l,~{}\forall b,$ are positive tolerances, and $N_{\mathcal{C}}(\mathcal{X})$ and $N_{\overline{\mathcal{C}}}(\overline{\mathcal{X}})$ are the general normal cone of $\mathcal{C}$ and $\overline{\mathcal{C}}$, at points $\mathcal{X}$ and $\overline{\mathcal{X}}$, respectively. On the other hand, the optimality conditions of the ADMM blocks 1 and 2, problems (29b) and (29c), result in:
00 0 $\in$ $\nabla f(\mathcal{X}^{(t)})+\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(t-1)}_{l,b}+\sum_{b\in\mathcal{B}}\Bigl{[}\rho^{(k)}_{s,b}\bigl{(}\mathbf{s}^{(t)}_{b}-\overline{\mathbf{s}}^{(t-1)}_{b}+\mathbf{z}^{(t-1)}_{s,b}\bigr{)}+\rho^{(k)}_{I,b}\bigl{(}\mathbf{I}^{(t)}_{b}-\overline{\mathbf{I}}^{(t-1)}_{b}+\mathbf{z}^{(t-1)}_{I,b}\bigr{)}\Bigr{]}+N_{\mathcal{C}}(\mathcal{X}^{(t)}),\qquad$ (37a)
00 0 $\in$ $-\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(t-1)}_{l,b}-\sum_{b\in\mathcal{B}}\Bigl{[}\rho^{(k)}_{s,b}\bigl{(}\mathbf{s}^{(t)}_{b}-\overline{\mathbf{s}}^{(t)}_{b}+\mathbf{z}^{(t-1)}_{s,b}\bigr{)}+\rho^{(k)}_{I,b}\bigl{(}\mathbf{I}^{(t)}_{b}-\overline{\mathbf{I}}^{(t)}_{b}+\mathbf{z}^{(t-1)}_{I,b}\bigr{)}\Bigr{]}+N_{\overline{\mathcal{C}}}(\overline{\mathcal{X}}^{(t)}).$ (37b)
Considering the Lagrange variable update of the inner algorithm, the optimality conditions of ADMM blocks 1 and 2 can be expressed as
$\sum_{b\in\mathcal{B}}\Bigl{[}\rho^{(k)}_{s,b}\bigl{(}-\overline{\mathbf{s}}^{(t)}_{b}+\mathbf{z}^{(t)}_{s,b}+\overline{\mathbf{s}}^{(t-1)}_{b}-\mathbf{z}^{(t-1)}_{s,b}\bigr{)}+\rho^{(k)}_{I,b}\bigl{(}-\overline{\mathbf{I}}^{(t)}_{b}+\mathbf{z}^{(t)}_{I,b}+\overline{\mathbf{I}}^{(t-1)}_{b}-\mathbf{z}^{(t-1)}_{I,b}\bigr{)}\Bigr{]}\in\nabla f(\mathcal{X}^{(t)})+\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(t)}_{l,b}+N_{\mathcal{C}}(\mathcal{X}^{(t)}),\qquad$ (38a)
$\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\Bigl{[}\rho^{(k)}_{l,b}\bigl{(}\mathbf{z}^{(t-1)}_{l,b}-\mathbf{z}^{(t)}_{l,b}\bigr{)}\Bigr{]}\in-\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(t)}_{l,b}+N_{\overline{\mathcal{C}}}(\overline{\mathcal{X}}^{(t)}).$ (39a)
Comparing (38a) and (36a), the stopping criteria for inner-level algorithm 2 will be according to (35a).
V-A2 Stopping criteria for outer-level Algorithm 3
The outer-level Algorithm 3 aims at implementing ALM to move $\mathbf{z}_{l,b},~{}\forall b,~{}\forall l$, to zero. Furthermore, upon convergence, each BS will consider the beamforming weight matrices $\mathbf{W}^{b},~{}\forall b,$ as the final decision. As a result, we consider the outer-level ALM algorithm to stop when the $\mathbf{z}_{l,b}$ values are small enough and the increase in the PF objective function value is less than a threshold, i.e.,:
$\left\|\mathbf{z}^{(k)}_{l,b}\right\|\leq\epsilon_{o,1},~{}\forall l,~{}\forall b,\qquad$ (40a)
$\frac{\big{|}f(\mathcal{X}^{(k)})-f(\mathcal{X}^{(k-1)})\big{|}}{\big{|}f(\mathcal{X}^{(k-1)})\big{|}}<\epsilon_{{o,2}}.$ (40b)
V-B Convergence
In order to discuss the convergence of the proposed algorithm, we make below three assumptions, which hold for the formulated problem (IV-A), considering sufficient available transmit power at each BS.
Assumption 1.
Problem (IV-A) is feasible, and the set of stationary points of problem (III) is nonempty.
Assumption 2.
The objective function $f(\mathcal{X})$ is continuously differentiable, $\mathcal{C}$ is a compact set, and $\overline{\mathcal{C}}$ is convex and compact.
Assumption 3.
Given $\boldsymbol{\lambda}^{(k)}_{l,b}$, $\rho^{(k)}_{o}$, and $\rho^{(k)}_{l,b},~{}\forall l,~{}\forall b$, the first block Algorithm 1 can find a stationary solution $\mathcal{X}^{(t)}$ such that
$0\in\partial_{x}L(\mathcal{X}^{(t)},\overline{\mathcal{X}}^{(t-1)},\mathcal{Z}^{(t-1)},\boldsymbol{\Psi}^{(t-1)}),$
and
$L(\mathcal{X}^{(t)},\overline{\mathcal{X}}^{(t-1)},\mathcal{Z}^{(t-1)},\boldsymbol{\Psi}^{(t-1)})\leq L(\mathcal{X}^{(t-1)},\overline{\mathcal{X}}^{(t-1)},\mathcal{Z}^{(t-1)},\boldsymbol{\Psi}^{(t-1)})<+\infty,$
for all $t\in\mathbb{Z}_{++}$.
V-B1 Convergence of the inner-level algorithm 2
According to [[32], Proposition 1], under the above assumptions, the inner-level Algorithm 2 terminates at a point $(\mathcal{X}^{(k)},\overline{\mathcal{X}}^{(k)},\mathcal{Z}^{(k)})$ satisfying the conditions (36a).
V-B2 Convergence of the outer-level algorithm 3
Under assumptions 1-3, according to [[32], Theorem 1], the outer-level Algorithm has a limit point $(\mathcal{X}^{*},\overline{\mathcal{X}}^{*},\mathcal{Z}^{*})$ which is either feasible to the original problem, i.e., $\mathcal{Z}^{*}=0,$ or $(\mathcal{X}^{*},\overline{\mathcal{X}}^{*})$ is a stationary point of the problem
$\underset{\mathcal{X}\in\mathcal{C},~{}\overline{\mathcal{X}}\in\overline{\mathcal{C}}}{\operatorname{minimize}}\,\,~{}\frac{1}{2}\left\|\mathcal{X}-\overline{\mathcal{X}}\right\|^{2}.$ (41)
According to [[32], Theorem 2], if $\boldsymbol{\Psi}^{*}$ is the limit Lagrange multiplier, associated with the point $(\mathcal{X}^{*},\overline{\mathcal{X}}^{*},\mathcal{Z}^{*})$, the point $(\mathcal{X}^{*},\overline{\mathcal{X}}^{*},\boldsymbol{\Psi}^{*})$ is a stationary point of the original problem (II-B).
V-C Computational Complexity and Signaling Overhead
In this subsection, we compare the per-iteration computational complexity of the centralized and distributed algorithms. For the distributed algorithm, we focus on the local optimization cost of the ADMM block 2, which is the dominant contributor to the per-iteration complexity. Specifically, ADMM block 2 subproblems take the form of SOC programs (SOCP) with constraints such as power limits and interference bounds that are SOC representable. The worst-case complexity of a primal-dual interior-point method for SOCP, as implemented in solvers such as mosek or SeDuMi [38, 39, 37], scales approximately as $\mathcal{O}(n^{3.5})$, where $n$ denotes the total cone dimension. For ADMM Block 2, each BS $b$ solves a local SOCP whose total cone dimension is approximately $2U^{2}+2N_{b}U$. Therefore, the per-iteration worst-case computational complexity at BS $b$ is approximately $\mathcal{O}\bigl{(}(U^{2}+N_{b}U)^{3.5}\bigr{)}.$ Considering HAPS with $N_{B+1}$ antenna elements, typically much larger than that of an MBS, the per-iteration complexity of the distributed approach is therefore upper bounded by $\mathcal{O}\bigl{(}(U^{2}+N_{B+1}U)^{3.5}\bigr{)}$. In contrast, the centralized algorithm jointly optimizes over all BSs. Its problem dimension aggregates all local variables, leading to a total cone dimension proportional to $\sum_{b\in\mathcal{B}}(U^{2}+N_{b}U)$. The resulting worst-case complexity per iteration is approximately $\mathcal{O}\Bigl{(}\bigl{(}\sum_{b\in\mathcal{B}}(U^{2}+N_{b}U)\bigr{)}^{3.5}\Bigr{)},$ which scales much worse in the network size, especially as the number of BSs or antenna elements grows. Furthermore, the distributed approach offers significant improvements in signaling overhead. In the centralized scheme, each BS must transmit all local channel state information to a central controller, amounting to $2N_{b}U$ real scalars per BS. By contrast, in the distributed scheme, only the consensus variables need to be exchanged, requiring approximately $3U$ real scalars per BS. This reduced overhead is particularly advantageous when $N_{b}$ is large, specifically in the HAPS case.
VI Numerical Result and Discussion
TABLE I: Simulation Parameters.
Parameter Value
HAPS altitude $20$ Km
Center frequency ($f_{c}$) $2.545$ GHz
[…]
$d_{x},~{}d_{y}$ $\lambda/2$
$\sigma^{2}_{n}$ $-100$ dBm
$N_{\text{b}},~{}\forall b\in\mathcal{B}\setminus\{B+1\},~{}N_{B+1}$ $4\times 4,~{}8\times 8$
$P_{b}^{\text{max}},~{}\forall b\in\mathcal{B}\setminus\{B+1\},~{}P_{B+1}^{\text{max}}$ $43$ dBm, $52$ dBm
Outer-level penalty parameter ($\rho^{(0)}_{o}$) $10$
$\delta$ $2$
$\omega$ $0.5$
$\gamma$ $1.5$
Figure 3: Convergence behavior of the proposed PFBWD algorithm.
In this section, we evaluate the performance of the proposed distributed PFBWD algorithm in HAPS-empowered vHetNets and compare it against both standalone terrestrial networks and the centralized approach. The standalone terrestrial network refers to a scenario where no HAPS is involved. In the case of a centralized algorithm, the original problem is solved using the SCA framework, which incorporates some RLT, as detailed in [4]. The simulated network consists of four MBSs and one HAPS, deployed over a $4~{}\text{km}\times 4~{}\text{km}$ square urban area. The HAPS and MBSs are equipped with $8\times 8$ and $4\times 4$ MIMO antenna arrays, respectively, and jointly serve $U$ single-antenna UEs in a cell-free architecture. The UEs are uniformly distributed across the considered area. The remaining simulation parameters and their values are summarized in Table I. All results are averaged over $1000$ independent network realizations. The algorithm is implemented using CVX parser, which runs on MATLAB, with mosek 9.1.9 as an internal solver [38].
TABLE II: Impact of inner-level penalty parameter on the required number of iterations.
Parameter $\delta$=0.5 $\delta$=1 $\delta$=2
Inner-level algorithm iterations $13.5$ $12.5$ $8.33$
(a) CDF of average SE per UE (b/s/Hz/UE).
(b) CDF of average PF objective function value.
Figure 4: Statistical behavior of SE per UE and the average PF objective function value in a HAPS-empowered vHetNet ($4$ MBSs + $1$ HAPS) serving $16$ UEs.
(a) Average SE per UE vs. number of UEs
(b) Average PF objective function value vs. number of UEs
Figure 5: Performance comparison of centralized and distributed PFBWD algorithms for different numbers of UEs in a HAPS-empowered vHetNet ($4$ MBSs + $1$ HAPS).
First, we discuss the convergence behavior of the proposed distributed PFBWD algorithm. Fig. 3 plots the average sum of residuals, defined as $\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\left\|\mathbf{r}_{l,b}\right\|$, over the inner-level ADMM Algorithm 2 iterations $t$, and the average PF objective function value (i.e., $\sum_{u\in\mathcal{U}}{\log(\log_{2}(1+\gamma_{u}))}$) over the outer-level Algorithm 3 iterations $k$. It can be observed that the residual values converge to zero as $t$ increases, while the objective value increases with $k$, demonstrating the convergence behavior of the proposed algorithm. The outer-level ALM algorithm also exhibits fast convergence as the objective value saturates after a few iterations (on an average $k=4$ iterations). However, the choice of penalty parameters plays a critical role in determining the number of required iterations. As shown in Table II, as the ratio of the inner-level to outer-level penalty parameters (i.e., $\delta$ in Algorithm 2) increases the number of required iterations reduces. This is because a higher penalty parameter enforces the satisfaction of the coupling constraints more aggressively, thereby accelerating convergence to feasibility. However, excessively large values of $\delta$ may lead to numerical instability [24].
Figs. 4a and  4b plot the cumulative distribution function (CDF) of the average SE per UE and the average PF objective value, respectively. The results compare the proposed distributed PFBWD algorithm with the centralized algorithm in a vHetNet comprising $4$ MBSs and $1$ HAPS, serving $16$ UEs. It can be observed that the distributed algorithm achieves lower average SE and objective function values. This is because the centralized algorithm has access to all network information, including complete channel coefficient matrices, enabling it to make decisions closer to the optimal solution. Conversely, in the distributed algorithm, each BS makes local decisions based on its channel information and coordinates with others by sharing selected variables. Consequently, the distributed approach yields slightly lower performance than the centralized counterpart. However, this performance gap is small, and the distributed algorithm’s results remain comparable to the centralized solution.
TABLE III: Comparison of computational complexity and signaling overhead for centralized vs. distributed PFBWD algorithms in a vHetNet ($4$ MBSs + $1$ HAPS) serving $16$ UEs.
Parameter Distributed Centralized
Computational complexity $\mathcal{O}(1280^{3.5})$ $\mathcal{O}(3328^{3.5})$
Signaling overhead (real scalars per BS) $\approx 48$ $\approx 512$ for MBS $\approx 2048$ for HAPS
Furthermore, Fig. 5, illustrates the performance gap between the centralized and distributed PFBWD algorithms with increasing number of UEs. Specifically, Figs. 5(a) and  5(b) plot the average SE per UE and the average PF objective function value, respectively, in a HAPS-empowered vHetNet consisting of $4$ MBSs and $1$ HAPS. As the number of UEs grows, the average SE per UE decreases mainly due to two reasons: i) the available resources are shared among more UEs and ii) the increased interference. Additionally, the performance gap between the centralized and distributed algorithms slightly widens with increasing UEs. This is because the amount of missing information at each BS in the distributed approach grows, leading to a minor decline in performance. Importantly, this minor decrease in performance comes with significant gains in algorithmic complexity and reduced signaling overhead among BSs. Table III summarizes the computational complexity and signaling overhead requirements for both algorithms when serving $16$ UEs, considering $4\times 4$ and $8\times 8$ antenna arrays at MBSs and HAPS, respectively. The proposed PFBWD Algorithm 3 offers much lower complexity and substantially less data exchange, resulting in a significant reduction in signaling overhead.
Figure 6: Average SE per UE: vHetNet vs. standalone terrestrial network across different numbers of MBSs.
To further examine the impact of vHetNet architecture and HAPS capabilities on network performance, Fig. 6 compares the performance of the HAPS-empowered vHetNet and the standalone terrestrial network serving $16$ UEs, under the proposed PFBWD algorithm. First, it can be observed that increasing the number of MBSs in both vHetNet and standalone terrestrial networks improves the average SE per UE, primarily due to the availability of more resources, including a higher number of antenna elements. Second, integrating HAPS into the terrestrial network architecture yields a noticeable enhancement in the average SE per UE. Specifically, for all values of $B$, integrating the terrestrial network with one HAPS, resulting in a total of $(B+1)$ BSs, consistently leads to improved SE for UEs. Furthermore, it can be observed that to match the performance of the $(4+1)$ vHetNet configuration ($4$ MBSs and $1$ HAPS), the standalone terrestrial network requires a total of $12$ MBSs. In other words, the vHetNet with $4$ MBSs and $1$ HAPS achieves the same average SE per UE as a standalone terrestrial network comprising $12$ MBSs, clearly demonstrating the efficiency of integrating HAPS. Importantly, these results also highlight the scalability of the proposed distributed algorithm, which can efficiently solve the beamforming design problem even for large numbers of MBSs in the network.
Figure 7: Impact of HAPS antenna architecture on the average SE per UE in a vHetNet ($4$ MBSs + $1$ HAPS) serving $16$ UEs.
In line with this observation, Fig. 7 presents the average SE per UE in a $(4+1)$ vHetNet configuration serving $16$ UEs, for various HAPS antenna configurations. As shown, the SE gain provided by HAPS integration can be further amplified by increasing the number of antenna elements at the HAPS. This is particularly promising, as the large surface area of HAPS platforms allows for accommodating a significantly larger number of antennas, enabling improved beamforming capabilities and higher spatial resolution.
VII Conclusion
This paper addressed the challenges of interference management in harmonized spectrum HAPS-empowered vertical heterogeneous networks (vHetNets). We highlighted the limitations of centralized interference management approaches, which face scalability issues due to the large number of MBSs, high-dimensional antenna arrays, and significant signaling overhead. To overcome these challenges, we formulated a cell-free proportional fairness beamforming weight design (PFBWD) problem tailored for HAPS-empowered vHetNet deployments. Recognizing the nonconvex nature of the problem, we adopted a two-level distributed algorithm that embeds a structured three-block ADMM within an augmented Lagrangian framework, ensuring guaranteed convergence. Simulation studies confirm convergence, effectiveness and scalability of the algorithm, showing that the algorithm can efficiently handle large-scale network configurations with many MBSs and high-dimensional antenna arrays. These findings underscore the practical value of advanced distributed optimization methods for enabling efficient and scalable interference management in future integrated aerial-terrestrial networks.